OpenStudy (anonymous):

Use implicit differentiation to find the slope of the tangent line to the curve -2 x^2 - 2 xy + 3 y^3 = -21 at the point ( 3,1 ). i'm so lost here what to do, please help

OpenStudy (jhannybean):

@perl cant type, im enjoying my feast >:')

OpenStudy (fibonaccichick666):

so, if you are differentiating with respect to y, what is the derivative of just y?

OpenStudy (fibonaccichick666):

same question, only you are differentiating with respect to x, what is the derivative of x?

OpenStudy (perl):

-2 x^2 - 2 xy + 3 y^3 = -21 use the product rule (and chain rule) on xy and chain rule on y^3

OpenStudy (fibonaccichick666):

@ericksol96

OpenStudy (perl):

-2x^2 - 2xy + 3y^3 = -21 the implicit derivative : -2*2*x - 2 ( 1*y + x* y ' ) + 3* 3* y^2 * y ' = 0

OpenStudy (perl):

any questions about that?

OpenStudy (nincompoop):

my question is why hasn't the person asking the question tried solving this himself first?

OpenStudy (swissgirl):

nin he wants some direction in life

OpenStudy (perl):

maybe technical difficulties

OpenStudy (nincompoop):

TRY solving this so we can help you help yourself.

OpenStudy (nincompoop):

HAPPY TURKEY night, Americans and US immigrants (illegal ones too)

OpenStudy (kainui):

@nincompoop can I give you two medals somehow?

OpenStudy (nincompoop):

thanks! I am stuffed, man